OpenStudy (anonymous):

(s-3 over 5 )^2) i need work show step by step

jimthompson5910 (jim_thompson5910):

\[\Large \left(\frac{s-3}{5}\right)^2\] \[\Large \frac{(s-3)^2}{5^2}\] \[\Large \frac{(s-3)(s-3)}{25}\] \[\Large \frac{s^2-3s-3s+9}{25}\] \[\Large \frac{s^2-6s+9}{25}\] So \[\Large \left(\frac{s-3}{5}\right)^2=\frac{s^2-6s+9}{25}\]